Eric is 2 times taller than Kevin but Kevin is 3 feet shorter than twice the height of Eric. What is each boy's height?
Answer by JoelSchwartz(130)

e=Eric's height in feet
k=Kevin's height in feet
e=2k
k=2e-3
e/2=k
e/2=2e-3
e/2-e/2+3=2e-3+3-e/2
3=1.5e
e=2
k=1